# A spin on Hagedorn temperatures and string stars

### v1

@@ 3.3 Grand canonical properties
In the second line, following Chen:2021dsw , we used the explicit $\beta$ dependence of the mass and the equations of motion to write the mass as a local integral. In the third line, we took the direct derivative of the action (75). One can check explicitly that the two agree by the equations of motion. Notice that since $2<d<6$, both terms in the third line are positive.
@@ 3.4 Microcanonical properties
The leading piece is just the Hagedorn behaviour. The other $J$-independent contribution is the leading correction due to gravitational self-interaction discussed by Chen:2021dsw , while the $J$-dependent piece is the leading correction due to angular momentum. Note that the latter is always negative: some of the energy has to be invested into (entropically suppressed) angular momentum.
@@
where $\tilde{l}_{0}$, $\tilde{c}_{||}$ and $\tilde{c}_{\perp}$ are numerical coefficients. Their values for $d=3,5$ are listed in Table 2.202020We don’t list $d=4$ as in that case $\beta-\beta_{H}$ is to leading order independent of $M$, which means that we need to take the next order into account to go to the microcanonical ensemble Chen:2021dsw . For the eccentricity (72) we find
@@ 4 The correspondence principle
The main importance of the string star saddle was the suggestion by Horowitz and Polchinski Horowitz:1997jc that it interpolates between the black hole and the free string phases of string theory. More precisely Chen:2021dsw , the suggestion is that as a Euclidean saddle defined close to $\beta\sim\beta_{H}$, it is continuously connected to the Euclidean black hole at lower temperatures $\beta\gg\beta_{H}$. In other words, as the Euclidean black hole size reaches the string scale, its Einstein gravity description is no longer valid, and should now be described as a self-gravitating string. In this context, this is what is usually referred to as the correspondence principle, or the black hole-string transition. The main argument in favor of this suggestion was that extrapolating the thermodynamic properties of both saddles to an intermediate regime $\beta-\beta_{H}\sim l_{s}$ (where both are ill defined), gives qualitatively the same result.
@@ 4.1 Rotating black holes
We begin by discussing the properties of slowly rotating black holes, following the conventions of Emparan:2003sy ; Ceplak:2023afb . The thermodynamic properties of rotating black holes in $D=d+1$ dimensions (also known as Myers-Perry black holes Myers:1986un ) are best described using the variables $\mu$, $a$, and the black hole horizon radius $r_{h}$. These variables are related by the condition
@@
To describe the shape of the black hole, we would like to define the black hole eccentricity $e=\sqrt{1-r_{\perp}^{2}/r_{||}^{2}}$. There are several ways to define the radii of a rotating black hole Ceplak:2024dxm . We make the following choice, following Ceplak:2024dxm : for the polar radius, we can measure the area of the horizon along the remaining d3d-3 directions and extract the radius, while for the equator radius, we measure the horizon area in the equatorial plane. This choice gives the radii
@@
Using the results of section 3.3, we can extrapolate the rotating string star properties (71), (72), (76), (77), (78) to the intermediate regime $\alpha^{\prime}m_{0}^{2}\sim 1$ (or $R^{2}-R_{H}^{2}\sim\alpha^{\prime}$Chen:2021dsw which gives
@@
First, note that for both saddles the perturbative parameter at the correspondence region is $l_{s}\Omega$. At zeroth order, they qualitatively match, as was already shown in the non-rotating case Chen:2021dsw . The angular momentum and eccentricity are the only quantities with a leading correction in $\Omega$. We see that in both cases, the values match up to $O(1)$ factors between the two saddles, which is the best we can hope for in this situation. We stress that the sign also matches, implying that both saddles rotate in the direction of the angular velocity $\Omega$, and become more oblate in the parallel plane due to the centrifugal force. [^21] Since at this order $J=\mathcal{I}\cdot\Omega$ on both sides, it also means that the moment of inertia qualitatively agrees between the two saddles.
@@ 4.2 Rotating strings
As the size of the string star increases, it becomes more diluted, and its internal forces decrease. From the other direction, as a string state turns heavier, its internal gravitational forces grow until they can no longer be neglected or treated perturbatively. We therefore expect a transition between the free string phase and the string star saddle. We will follow section 2 and describe the properties of both sides in the microcanonical ensemble. [^22] The canonical single-string ensemble suffers from multiple subtleties such as negative specific heat Mertens:2015ola . To estimate when the gravitational forces of the string become strong, we can compare the gravitational potential to the free string energy level separation
@@
At subleading order, the corrections to the sizes generally have a different sign. Similarly to the black hole transition region, this doesn’t necessarily indicate a problem and depends on the $J$ dependence of the breakdown region. One way to define the breakdown region is by demanding the string star’s on-shell action to be $O(1)$ (some $J$-independent constant) Chen:2021dsw . In this way, it is possible to show that the sizes of both the free strings and the string star get a positive correction in $J^{2}$ at the breakdown scale. With better control over the proper correspondence region, one could hope for a match beyond the perturbative expansions. [^24] For the regime $(R-R_{H}(\nu))/l_{s}\sim|\nu|$ mentioned in section 3.2, the breakdown of the semiclassical solution happens not at a specific temperature, but for $|\nu|\lesssim g_{s}^{\frac{4}{6-d}}$. At that point, $M\sim g_{s}^{\frac{4}{d-6}}$, $\mathcal{I}\sim M^{2}$, $J\sim iM$, so one might speculate that this phase is connected to free strings with $J\sim M$. This regime is not connected to the black hole phase due to the condition $\beta-\beta_{H}\lesssim|\nu|\ll 1$.
A more demanding test for the correspondence would be to analyze the situation at $J\sim g_{s}^{\frac{4}{d-6}}$, or $J\gg g_{s}^{\frac{4}{d-6}}$, corresponding to $J\sim M$ or $J\gg M$ in the free string description. The correspondence point should then be affected at leading order in $J$. Indeed, for $J\sim M^{2}$, we are in the ultraspinning regime, which has been analyzed in Ceplak:2023afb ; Ceplak:2024dxm . Again, this requires more control over the string star solution than we currently have.
@@ 4.3 Large angular momentum and extremal objects
It is tempting to speculate about the string star solution closer to the string Regge bound $J\lesssim\frac{1}{2}\alpha^{\prime}M^{2}$ (4Ceplak:2023afb ; Ceplak:2024dxm . Lacking control over such a large $\Omega$ regime, it is hard to predict the result. A natural suggestion would be that in this regime, the string star is composed of near-extremal strings, with small $O(1)$ entropy. In this picture, one could expect the eccentricity to be very close to unity, with $L_{\perp}\sim l_{s}$. This ‘stringy pancake’ limit is likely to transition into a ‘quantum string bar’. It would be interesting to better understand this regime.
Finally, one could wonder about the relation between our solution to extremal rotating black holes, which also satisfy $J\propto M^{2}$ (albeit with a different $g_{s}$ factor Ceplak:2023afb ). The extremal limit necessarily involves arbitrarily low temperatures and a very large horizon area. As a result, we believe that in order to potentially deform an (Eudlidean) extremal black hole to a string star, one has to first go through higher temperatures $\beta\sim l_{s}$.
@@ 5 Anti-de Sitter space
with $J$ being a rotation generator in $S^{d-1}$. [^25] Notice that here both $\beta$ and $\Omega$ are defined on the CFT side in dimensionless units. In terms of the dual string theory, they are in $l_{ads}$ units. As a result, they are related to the previous section’s $\beta$, $\Omega$ by factors of $l_{ads}$. This ensemble is (holographically dual to) the AdS analog of (1). Unlike in flat space, where one fixes the asymptotic length of the thermal circle, in AdS one fixes the asymptotic form of the metric (or equivalently the conformal boundary)Hawking:1982dh ; Witten:1998qj . This prescription renders the AdS gravitational partition function well defined even at finite string coupling, in contrast to the situation in flat space.
@@ 5.1 The Hagedorn temperature of rotating thermal AdS
As was shown in maldacena_private ; Urbach:2022xzw , the leading correction in $l_{s}/l_{ads}$ of the Hagedorn temperature can be computed using the effective action (40), adapted to curved space
@@
where in the second line we expanded to quadratic order in $r$. We already see a dramatic difference compared to flat space. Due to the AdS curvature, the winding mode remains stable even with a real chemical potential, as long as $\Omega<1$. The instability of the ensemble for $\Omega\geq 1$ is a direct result of the unitarity bound $E\geq J$ (in AdS units) Hawking:1998kw ; Hawking:1999dp ; Witten:2021nzp ; Kim:2023sig .
@@
Following the Newtonian intuition, as we increase the angular velocity $\Omega$, the solution becomes wider in the 2d plane. When $1-\Omega^{2}\sim l_{s}/l_{ads}$, we expect higher order kinetic terms to be leading in the plane, eventually leading to an $L_{||}\sim l_{ads}$ sized solution Urbach:2023npi .
@@
As we can see, the solution size in the 2d plane is significantly smaller compared to the perpendicular direction $L_{||}\ll L_{\perp}$. The intuition is that in the 2d plane, the solution is effectively the flat space solution, giving the first term in (110). In the other d2d-2 directions, the solution is the standard AdS result maldacena_private ; Urbach:2022xzw , which leads to the second term in (110). It would be interesting to check how higher-order corrections depend on $\nu$.
@@
In AdS3 ($d=2$), we can consider backgrounds with non-zero NS-NS flux $\lambda$ Cho:2018nfn ; Urbach:2023npi . To leading $\alpha^{\prime}$ order, the effect of $\lambda$ on the quadratic action of the winding mode is the same as $\Omega$. Shifting $\Omega^{2}\mapsto\Omega^{2}+\lambda^{2}$ in (109) gives the leading correction to the Hagedorn temperature
@@
Notice that this time, the ensemble is unstable for $\Omega^{2}+\lambda^{2}\geq 1$ due to condensation of rotating long-strings at large radii. Following Harmark:2024ioq ; Canneti:2025cos , it would be interesting to study higher order corrections to the Hagedorn temperature at order $1/l_{AdS}^{2}$, $1/l_{AdS}^{3}$ as a function of the NS-NS flux $\lambda$ and angular momentum $\Omega$. In Harmark:2024ioq , the higher $\alpha^{\prime}$-corrections to the winding mode equation were found by reverse-engineering from the exactly known Hagedorn temperature on the pp-wave background. As they are stated in covariant form, one can also use these expressions to find the $\alpha^{\prime}$-corrections to the Hagedorn temperature in the ensemble with nonzero $\nu$.
The analysis above breaks for the pure NS-NS background $\lambda=1$. At $\Omega=0$, the magnetic flux and the gravitational contributions to the effective mass cancel. As a result, the winding mode has no potential and extends with $L\sim l_{ads}$ Urbach:2023npi . Fortunately, the pure NS-NS background has an RNS description, which allows for a direct calculation of the spectrum (see the recent Ferko:2024uxi ). The string spectrum of bosonic strings in thermal pure NS-NS AdS3 for general $\beta,\nu$ was studied in Mertens:2014nca , where the authors also found the Hagedorn temperature $R_{H}(\nu)$. Modifying their computation for type II strings, we find ($1<\nu<1$)
@@
with $k=l_{ads}^{2}/\alpha^{\prime}$. Surprisingly, this temperature agrees exactly with the flat space result (34) for $k|\nu|>1$. Thus, it also agrees for any $\nu$ in the flat space limit $k\gg 1$. It also converges to the known result at $\nu=0$ Berkooz:2007fe . The (second-derivative) discontinuity of $R^{2}_{H}$ at $k|\nu|=1$ is the location where the $SL(2,R)$ representation of the winding mode turns from continuous at $k|\nu|<1$ to discrete at $k|\nu|>1$. As we increase $\nu$, the size of the winding mode gets shorter. We interpret it as the transition where generic strings turn from long strings to short strings Maldacena:2000hw . It would be interesting to derive (113) directly using the recent results of Ferko:2024uxi .
@@ 5.2 Rotating AdS string stars
In Urbach:2022xzw , a string star saddle was found numerically in AdS space. The solution size was shown to grow with the temperature, with a maximal size of $L^{2}\sim l_{s}l_{ads}$ at the AdS Hagedorn temperature. At lower temperatures $l_{s}/l_{ads}\ll(R-R_{H})/R_{H}\ll 1$ (and $3\leq d\leq 5$), the star’s size is much smaller than the AdS radius, and thus approximately given by the flat space string star configuration. Similarly, the (small) AdS black hole at high temperatures $l_{s}\ll\beta\ll l_{ads}$ is well approximated by the Schwarzschild solution. As a result, the qualitative arguments for the string/black hole correspondence at intermediate temperatures $R-R_{H}\sim l_{s}$ in flat space immediately apply to AdS.
@@
Expanding in small $\nu\ll 1$, we expect a perturbation of the non-rotating AdS string star. To leading non-trivial order, we can use the rigid-body approximation (80). Using Urbach:2022xzw , we can estimate the moment of inertia close to the AdS Hagedorn temperature $(R-R_{H})/R_{H}\ll l_{s}/l_{ads}\ll 1$:
@@
Near the AdS Hagedorn temperature, we also expect the solution to have a size similar to that of the winding mode described in the previous section. Note that due to the AdS curvature attraction, it is possible in this case to find real $\Omega$ solutions for the string star directly, without the need to go to imaginary angular velocity. At lower temperatures, however, we expect the size to shrink. For $l_{s}/l_{ads}\ll(R-R_{H})/R_{H}\ll 1$ and $3\leq d\leq 5$ (and real $\nu$), the solution coincides with the flat-space rotating string star solution of section 3. As in $\nu=0$, we thus expect the arguments of section 4 to follow to AdS as well. [^26] For recent analyses of the superradiant instability of rotating AdS black holes and the proposed end-state in the form of ‘grey galaxies’ (mixed configurations of a near-critical $\Omega\sim 1$ black hole and a co-rotating thermal cloud) see Kim:2023sig ; Bajaj:2024utv .
A nuanced version of this discussion exists for AdS3. While in $d=2$ flat space no string star solutions exist, it was shown in Urbach:2023npi that such solutions do exist in AdS3. Following Halder:2022ykw , it was suggested that the AdS3 string star is continuously related to the ‘S-transformed’ BTZ solution deformed by an angular winding condensate (defined only for very low temperatures). Adding non-zero angular velocity, we expect a generalization of this discussion. In terms of the modular parameter $\tau=\nu+i\beta/(2\pi)$, we can define the Hagedorn line $\tau_{H}=\nu+iR_{H}(\nu)/l_{ads}$. A reliable string star solution exists in a small region close to $\tau_{H}$. [^27] More precisely, $\tau=\nu+iR/l_{ads}$ is in the region if $\nu\ll 1$ and $R-R_{H}(\nu)\ll l_{s}$. Using the S-transformation $\tau^{\prime}=-1/\tau$, there also is a second region close to $1/\tau_{H}$ where we can find a reliable winding condensate, this time around the (rotating) BTZ solution. The string/BTZ transition suggested in  Halder:2022ykw ; Urbach:2023npi can be understood as the claim that these two regions are smoothly related as a function of $\tau$.

### v2

@@ 3.3 Grand canonical properties
In the second line, following Chen et al. (2023), we used the explicit $\beta$ dependence of the mass and the equations of motion to write the mass as a local integral. In the third line, we took the direct derivative of the action (75). One can check explicitly that the two agree by the equations of motion. Notice that since $2<d<6$, both terms in the third line are positive.
@@ 3.4 Microcanonical properties
The leading piece is just the Hagedorn behaviour. The other $J$-independent contribution is the leading correction due to gravitational self-interaction discussed by Chen et al. (2023), while the $J$-dependent piece is the leading correction due to angular momentum. Note that the latter is always negative: some of the energy has to be invested into (entropically suppressed) angular momentum.
@@
where $\tilde{l}_{0}$, $\tilde{c}_{||}$ and $\tilde{c}_{\perp}$ are numerical coefficients. Their values for $d=3,5$ are listed in Table 2.202020We don’t list $d=4$ as in that case $\beta-\beta_{H}$ is to leading order independent of $M$, which means that we need to take the next order into account to go to the microcanonical ensemble Chen et al. (2023). For the eccentricity (72) we find
@@ 4 The correspondence principle
The main importance of the string star saddle was the suggestion by Horowitz and Polchinski Horowitz and Polchinski (1998) that it interpolates between the black hole and the free string phases of string theory. More precisely Chen et al. (2023), the suggestion is that as a Euclidean saddle defined close to $\beta\sim\beta_{H}$, it is continuously connected to the Euclidean black hole at lower temperatures $\beta\gg\beta_{H}$. In other words, as the Euclidean black hole size reaches the string scale, its Einstein gravity description is no longer valid, and should now be described as a self-gravitating string. In this context, this is what is usually referred to as the correspondence principle, or the black hole-string transition. The main argument in favor of this suggestion was that extrapolating the thermodynamic properties of both saddles to an intermediate regime $\beta-\beta_{H}\sim l_{s}$ (where both are ill defined), gives qualitatively the same result.
@@ 4.1 Rotating black holes
We begin by discussing the properties of slowly rotating black holes, following the conventions of Emparan and Myers (2003); Čeplak et al. (2023). The thermodynamic properties of rotating black holes in $D=d+1$ dimensions (also known as Myers-Perry black holes Myers and Perry (1986)) are best described using the variables $\mu$, $a$, and the black hole horizon radius $r_{h}$. These variables are related by the condition
@@
To describe the shape of the black hole, we would like to define the black hole eccentricity $e=\sqrt{1-r_{\perp}^{2}/r_{||}^{2}}$. There are several ways to define the radii of a rotating black hole Čeplak et al. (2024). We make the following choice, following Čeplak et al. (2024): for the polar radius, we can measure the area of the horizon along the remaining d3d-3 directions and extract the radius, while for the equator radius, we measure the horizon area in the equatorial plane. This choice gives the radii
@@
Using the results of section 3.3, we can extrapolate the rotating string star properties (71), (72), (76), (77), (78) to the intermediate regime $\alpha^{\prime}m_{0}^{2}\sim 1$ (or $R^{2}-R_{H}^{2}\sim\alpha^{\prime}$Chen et al. (2023) which gives
@@
First, note that for both saddles the perturbative parameter at the correspondence region is $l_{s}\Omega$. At zeroth order, they qualitatively match, as was already shown in the non-rotating case Chen et al. (2023). The angular momentum and eccentricity are the only quantities with a leading correction in $\Omega$. We see that in both cases, the values match up to $O(1)$ factors between the two saddles, which is the best we can hope for in this situation. We stress that the sign also matches, implying that both saddles rotate in the direction of the angular velocity $\Omega$, and become more oblate in the parallel plane due to the centrifugal force. [^21] Since at this order $J=\mathcal{I}\cdot\Omega$ on both sides, it also means that the moment of inertia qualitatively agrees between the two saddles.
@@ 4.2 Rotating strings
As the size of the string star increases, it becomes more diluted, and its internal forces decrease. From the other direction, as a string state turns heavier, its internal gravitational forces grow until they can no longer be neglected or treated perturbatively. We therefore expect a transition between the free string phase and the string star saddle. We will follow section 2 and describe the properties of both sides in the microcanonical ensemble. [^22] The canonical single-string ensemble suffers from multiple subtleties such as negative specific heat Mertens (2015). To estimate when the gravitational forces of the string become strong, we can compare the gravitational potential to the free string energy level separation
@@
At subleading order, the corrections to the sizes generally have a different sign. Similarly to the black hole transition region, this doesn’t necessarily indicate a problem and depends on the $J$ dependence of the breakdown region. One way to define the breakdown region is by demanding the string star’s on-shell action to be $O(1)$ (some $J$-independent constant) Chen et al. (2023). In this way, it is possible to show that the sizes of both the free strings and the string star get a positive correction in $J^{2}$ at the breakdown scale. With better control over the proper correspondence region, one could hope for a match beyond the perturbative expansions. [^24] For the regime $(R-R_{H}(\nu))/l_{s}\sim|\nu|$ mentioned in section 3.2, the breakdown of the semiclassical solution happens not at a specific temperature, but for $|\nu|\lesssim g_{s}^{\frac{4}{6-d}}$. At that point, $M\sim g_{s}^{\frac{4}{d-6}}$, $\mathcal{I}\sim M^{2}$, $J\sim iM$, so one might speculate that this phase is connected to free strings with $J\sim M$. This regime is not connected to the black hole phase due to the condition $\beta-\beta_{H}\lesssim|\nu|\ll 1$.
A more demanding test for the correspondence would be to analyze the situation at $J\sim g_{s}^{\frac{4}{d-6}}$, or $J\gg g_{s}^{\frac{4}{d-6}}$, corresponding to $J\sim M$ or $J\gg M$ in the free string description. The correspondence point should then be affected at leading order in $J$. Indeed, for $J\sim M^{2}$, we are in the ultraspinning regime, which has been analyzed in Čeplak et al. (2023, 2024). Again, this requires more control over the string star solution than we currently have.
@@ 4.3 Large angular momentum and extremal objects
It is tempting to speculate about the string star solution closer to the string Regge bound $J\lesssim\frac{1}{2}\alpha^{\prime}M^{2}$ (4Čeplak et al. (2023, 2024). Lacking control over such a large $\Omega$ regime, it is hard to predict the result. A natural suggestion would be that in this regime, the string star is composed of near-extremal strings, with small $O(1)$ entropy. In this picture, one could expect the eccentricity to be very close to unity, with $L_{\perp}\sim l_{s}$. This ‘stringy pancake’ limit is likely to transition into a ‘quantum string bar’. It would be interesting to better understand this regime.
Finally, one could wonder about the relation between our solution to extremal rotating black holes, which also satisfy $J\propto M^{2}$ (albeit with a different $g_{s}$ factor Čeplak et al. (2023)). The extremal limit necessarily involves arbitrarily low temperatures and a very large horizon area. As a result, we believe that in order to potentially deform an (Eudlidean) extremal black hole to a string star, one has to first go through higher temperatures $\beta\sim l_{s}$.
@@ 5 Anti-de Sitter space
with $J$ being a rotation generator in $S^{d-1}$. [^25] Notice that here both $\beta$ and $\Omega$ are defined on the CFT side in dimensionless units. In terms of the dual string theory, they are in $l_{ads}$ units. As a result, they are related to the previous section’s $\beta$, $\Omega$ by factors of $l_{ads}$. This ensemble is (holographically dual to) the AdS analog of (1). Unlike in flat space, where one fixes the asymptotic length of the thermal circle, in AdS one fixes the asymptotic form of the metric (or equivalently the conformal boundary)Hawking and Page (1983); Witten (1998). This prescription renders the AdS gravitational partition function well defined even at finite string coupling, in contrast to the situation in flat space.
@@ 5.1 The Hagedorn temperature of rotating thermal AdS
As was shown in Maldacena (2021); Urbach (2022), the leading correction in $l_{s}/l_{ads}$ of the Hagedorn temperature can be computed using the effective action (40), adapted to curved space
@@
where in the second line we expanded to quadratic order in $r$. We already see a dramatic difference compared to flat space. Due to the AdS curvature, the winding mode remains stable even with a real chemical potential, as long as $\Omega<1$. The instability of the ensemble for $\Omega\geq 1$ is a direct result of the unitarity bound $E\geq J$ (in AdS units) Hawking et al. (1999); Hawking and Reall (2000); Witten (2021); Kim et al. (2023).
@@
Following the Newtonian intuition, as we increase the angular velocity $\Omega$, the solution becomes wider in the 2d plane. When $1-\Omega^{2}\sim l_{s}/l_{ads}$, we expect higher order kinetic terms to be leading in the plane, eventually leading to an $L_{||}\sim l_{ads}$ sized solution Urbach (2023).
@@
As we can see, the solution size in the 2d plane is significantly smaller compared to the perpendicular direction $L_{||}\ll L_{\perp}$. The intuition is that in the 2d plane, the solution is effectively the flat space solution, giving the first term in (110). In the other d2d-2 directions, the solution is the standard AdS result Maldacena (2021); Urbach (2022), which leads to the second term in (110). It would be interesting to check how higher-order corrections depend on $\nu$.
@@
In AdS3 ($d=2$), we can consider backgrounds with non-zero NS-NS flux $\lambda$ Cho et al. (2020); Urbach (2023). To leading $\alpha^{\prime}$ order, the effect of $\lambda$ on the quadratic action of the winding mode is the same as $\Omega$. Shifting $\Omega^{2}\mapsto\Omega^{2}+\lambda^{2}$ in (109) gives the leading correction to the Hagedorn temperature
@@
Notice that this time, the ensemble is unstable for $\Omega^{2}+\lambda^{2}\geq 1$ due to condensation of rotating long-strings at large radii. Following Harmark (2024); Canneti (2025), it would be interesting to study higher order corrections to the Hagedorn temperature at order $1/l_{AdS}^{2}$, $1/l_{AdS}^{3}$ as a function of the NS-NS flux $\lambda$ and angular momentum $\Omega$. In Harmark (2024), the higher $\alpha^{\prime}$-corrections to the winding mode equation were found by reverse-engineering from the exactly known Hagedorn temperature on the pp-wave background. As they are stated in covariant form, one can also use these expressions to find the $\alpha^{\prime}$-corrections to the Hagedorn temperature in the ensemble with nonzero $\nu$.
The analysis above breaks for the pure NS-NS background $\lambda=1$. At $\Omega=0$, the magnetic flux and the gravitational contributions to the effective mass cancel. As a result, the winding mode has no potential and extends with $L\sim l_{ads}$ Urbach (2023). Fortunately, the pure NS-NS background has an RNS description, which allows for a direct calculation of the spectrum (see the recent Ferko et al. (2025)). The string spectrum of bosonic strings in thermal pure NS-NS AdS3 for general $\beta,\nu$ was studied in Mertens et al. (2014), where the authors also found the Hagedorn temperature $R_{H}(\nu)$. Modifying their computation for type II strings, we find ($1<\nu<1$)
@@
with $k=l_{ads}^{2}/\alpha^{\prime}$. Surprisingly, this temperature agrees exactly with the flat space result (34) for $k|\nu|>1$. Thus, it also agrees for any $\nu$ in the flat space limit $k\gg 1$. It also converges to the known result at $\nu=0$ Berkooz et al. (2007). The (second-derivative) discontinuity of $R^{2}_{H}$ at $k|\nu|=1$ is the location where the $SL(2,R)$ representation of the winding mode turns from continuous at $k|\nu|<1$ to discrete at $k|\nu|>1$. As we increase $\nu$, the size of the winding mode gets shorter. We interpret it as the transition where generic strings turn from long strings to short strings Maldacena and Ooguri (2001). It would be interesting to derive (113) directly using the recent results of Ferko et al. (2025).
@@ 5.2 Rotating AdS string stars
In Urbach (2022), a string star saddle was found numerically in AdS space. The solution size was shown to grow with the temperature, with a maximal size of $L^{2}\sim l_{s}l_{ads}$ at the AdS Hagedorn temperature. At lower temperatures $l_{s}/l_{ads}\ll(R-R_{H})/R_{H}\ll 1$ (and $3\leq d\leq 5$), the star’s size is much smaller than the AdS radius, and thus approximately given by the flat space string star configuration. Similarly, the (small) AdS black hole at high temperatures $l_{s}\ll\beta\ll l_{ads}$ is well approximated by the Schwarzschild solution. As a result, the qualitative arguments for the string/black hole correspondence at intermediate temperatures $R-R_{H}\sim l_{s}$ in flat space immediately apply to AdS.
@@
Expanding in small $\nu\ll 1$, we expect a perturbation of the non-rotating AdS string star. To leading non-trivial order, we can use the rigid-body approximation (80). Using Urbach (2022), we can estimate the moment of inertia close to the AdS Hagedorn temperature $(R-R_{H})/R_{H}\ll l_{s}/l_{ads}\ll 1$:
@@
Near the AdS Hagedorn temperature, we also expect the solution to have a size similar to that of the winding mode described in the previous section. Note that due to the AdS curvature attraction, it is possible in this case to find real $\Omega$ solutions for the string star directly, without the need to go to imaginary angular velocity. At lower temperatures, however, we expect the size to shrink. For $l_{s}/l_{ads}\ll(R-R_{H})/R_{H}\ll 1$ and $3\leq d\leq 5$ (and real $\nu$), the solution coincides with the flat-space rotating string star solution of section 3. As in $\nu=0$, we thus expect the arguments of section 4 to follow to AdS as well. [^26] For recent analyses of the superradiant instability of rotating AdS black holes and the proposed end-state in the form of ‘grey galaxies’ (mixed configurations of a near-critical $\Omega\sim 1$ black hole and a co-rotating thermal cloud) see Kim et al. (2023); Bajaj et al. (2024).
A nuanced version of this discussion exists for AdS3. While in $d=2$ flat space no string star solutions exist, it was shown in Urbach (2023) that such solutions do exist in AdS3. Following Halder et al. (2023), it was suggested that the AdS3 string star is continuously related to the ‘S-transformed’ BTZ solution deformed by an angular winding condensate (defined only for very low temperatures). Adding non-zero angular velocity, we expect a generalization of this discussion. In terms of the modular parameter $\tau=\nu+i\beta/(2\pi)$, we can define the Hagedorn line $\tau_{H}=\nu+iR_{H}(\nu)/l_{ads}$. A reliable string star solution exists in a small region close to $\tau_{H}$. [^27] More precisely, $\tau=\nu+iR/l_{ads}$ is in the region if $\nu\ll 1$ and $R-R_{H}(\nu)\ll l_{s}$. Using the S-transformation $\tau^{\prime}=-1/\tau$, there also is a second region close to $1/\tau_{H}$ where we can find a reliable winding condensate, this time around the (rotating) BTZ solution. The string/BTZ transition suggested in  Halder et al. (2023); Urbach (2023) can be understood as the claim that these two regions are smoothly related as a function of $\tau$.